# RDEIC: Accelerating Diffusion-Based Extreme Image Compression with Relay Residual Diffusion

### v2

@@ Abstract
Diffusion-based extreme image compression methods have achieved impressive performance at extremely low bitrates. However, constrained by the iterative denoising process that starts from pure noise, these methods are limited in both fidelity and efficiency. To address these two issues, we present Relay Residual Diffusion Extreme Image Compression (RDEIC), which leverages compressed feature initialization and residual diffusion. Specifically, we first use the compressed latent features of the image with added noise, instead of pure noise, as the starting point to eliminate the unnecessary initial stages of the denoising process. Second, we directly derive a novel residual diffusion equation from Stable Diffusion’s original diffusion equation that reconstructs the raw image by iteratively removing the added noise and the residual between the compressed and target latent features. In this way, we effectively combine the efficiency of residual diffusion with the powerful generative capability of Stable Diffusion. Third, we propose a fixed-step fine-tuning strategy to eliminate the discrepancy between the training and inference phases, thereby further improving the reconstruction quality. Extensive experiments demonstrate that the proposed RDEIC achieves state-of-the-art visual quality and outperforms existing diffusion-based extreme image compression methods in both fidelity and efficiency. The source code and pre-trained models will be released.
@@ I Introduction
In this work, we propose Relay Residual Diffusion Extreme Image Compression (RDEIC) to overcome the three limitations mentioned above. To overcome the first two limitations, we proposed a novel relay residual diffusion framework. Specifically, we construct the starting point using the compressed latent features combined with slight noise, transitioning between the starting point and target latent features by shifting the residual between them. This approach significantly reduces the number of denoising steps required for reconstruction while ensures that the starting point retains most of the information from the compressed features, providing a strong foundation for subsequent detail generation. To leverage the robust generative capability of pre-trained stable diffusion for extreme image compression, we derive a novel residual diffusion equation directly from stable diffusion’s diffusion equation, rather than designing a new diffusion equation from scratch as done in [17]. To address the third limitation, we introduce a fixed-step fine-tuning strategy to eliminate the discrepancy between the training and inference phases. By fine-tuning RDEIC throughout the entire reconstruction process, we further improve the reconstruction quality. Moreover, to meet users’ diverse requirements, we introduce a controllable detail generation method that achieves a trade-off between smoothness and sharpness by adjusting the intensity of high-frequency components in the reconstructions. As shown in Fig. 1, the proposed RDEIC achieves state-of-the-art perceptual performance at extremely low bitrates, and significantly outperforms existing diffusion-based extreme image compression methods with fewer inference steps.
@@
1) We propose a relay residual diffusion process that effectively combine the efficiency of residual diffusion with the powerful generative capability of stable diffusion. To the best of our knowledge, we are the first to successfully integrate stable diffusion into a residual diffusion framework.
2) To eliminate the discrepancy between the training and inference phases, we design a fixed-step fine-tuning strategy that refines the model through the entire reconstruction process, further improving reconstruction quality.
@@ II-A Learned Image Compression
As a pioneer work, Ballé et al. [18] proposed an end-to-end image compression framework to jointly optimize the rate-distortion performance. In [19], they later introduced a hyperprior to reduce spatial dependencies in the latent representation, greatly enhancing performance. Subsequent works further improved compression models by developing various nonlinear transforms [4, 6, 7, 20] and entropy models [21, 22, 23, 24]. However, optimization for rate-distortion alone often results in unrealistic reconstructions at low bitrates, typically manifested as blurring or over-smoothness [25].
To improve perceptual quality, generative models have been integrated into compression methods. Agustsson et al. [8] added an adversarial loss for lost details generation. Mentzer et al. [9] explored the generator and discriminator architectures, as well as training strategies for perceptual image compression. In [10], Muckley et al. introduced a local adversarial discriminator to enhance statistical fidelity. With the advancement of diffusion models, some efforts have been made to apply diffusion models to image compression. For instance, Yang et al. [11] innovatively introduced a conditional diffusion model as decoder for image compression. Kuang et al. [26] proposed a consistency guidance architecture to guide the diffusion model in stably reconstructing high-quality images.
@@ II-C Relay Diffusion
Conventional diffusion models, such as denoising diffusion probabilistic models (DDPM) [13] and its variants, operate by progressively adding and removing Gaussian noise through a multi-step Markov process. These models have achieved remarkable results in low-resolution scenarios but face substantial challenges in terms of computational efficiency and performance when applied to higher resolutions. To overcome this, cascaded diffusion methods [34, 35] decompose the image generation into multiple stages, with each stage responsible for super-resolution conditioning on the previous one. However, these methods still require complete resampling at each stage, leading to inefficiencies and potential mismatches among different resolutions.
Relay diffusion, as proposed by Teng et al. [36], extends the cascaded framework by continuing the diffusion process directly from the low-resolution output rather than restarting from pure noise, which allows the higher-resolution stages to correct artifacts from earlier stages. This design is particularly well-suited for tasks such as image restoration and image compression, where degraded images or features are available. For instance, PASD [37] and SeeSR [38] directly embed the LR latent into the initial random noise during the inference process to alleviate the inconsistency between training and inference. ResShift [17] further constructs a Markov chain that transfers between degraded and target features by shifting the residual between them, substantially improving the transition efficiency. However, its redesigned diffusion equation and noise schedule prevent it from leveraging the robust generative capability of pre-trained stable diffusion.
@@ III Methodology
Figure 2: The proposed RDEIC. We first map a raw image $\boldsymbol{x}$ into the latent space using the encoder $\mathcal{E}$ and then perform end-to-end lossy compression to get compressed latent features $\boldsymbol{z}_{c}$. We then use $\boldsymbol{z}_{c}$ with added noise as the starting point and apply a denoising process to reconstruct the noise-free latent feature $\boldsymbol{z}_{0}$. The decoder $\mathcal{D}$ maps $\boldsymbol{z}_{0}$ back to the pixel space, to get the reconstructed image $\hat{\boldsymbol{x}}$. (a) Vanilla diffusion framework that starts from pure noise. (b) The proposed relay residual diffusion framework that starts from compressed latent features with added noise.
@@ III-A Overall Framework
Fig. 2 shows an overview of the proposed RDEIC network. We first use an encoder $\mathcal{E}$ and analysis transform $g_{a}$ to convert the input image $\boldsymbol{x}$ to its latent representation $\boldsymbol{y}$. Then we perform hyper transform coding on $\boldsymbol{y}$ with the categorical hyper model [39] and use the space-channel context model $\mathcal{C}_{m}$ to predict the entropy parameters $(\boldsymbol{\mu},\boldsymbol{\sigma})$ to estimate the distribution of quantized latent representation $\boldsymbol{\hat{y}}$ [20]. The side information $\boldsymbol{l}_{p}$ is quantized through vector-quantization, i.e., $\hat{\boldsymbol{l}}_{p}$ is the mapping of $\boldsymbol{l}_{p}$ to its closest codebook entry. Subsequently, the synthesis transform $g_{s}$ is used to obtain the image content dependent features $\boldsymbol{z}_{c}$. Random noise is then added to $\boldsymbol{z}_{c}$, which is the starting point for reconstructing the noise-free latent features $\boldsymbol{z}_{0}$ through an iterative denoising process. The denoising process is implemented by a frozen pre-trained noise estimator $\epsilon_{sd}$ of stable diffusion with trainable control network for intermediate feature modulation. Finally, the reconstructed image $\boldsymbol{\hat{x}}$ is decoded from $\boldsymbol{z}_{0}$ using the decoder $\mathcal{D}$.
@@ III-B Relay Residual Diffusion
where $\epsilon_{t}\sim\mathcal{N}(0,\boldsymbol{I})$, $\alpha_{t}=1-\beta_{t}$ and $\bar{\alpha}_{t}=\prod_{i=1}^{t}\alpha_{i}$. When $t$ is large enough, the noisy latent $\boldsymbol{z}_{t}$ is nearly a standard Gaussian distribution. In practice, $T$ is typically very large, e.g., 1000, and pure noise is set as the starting point for the reverse diffusion process. However, this approach is not optimal for the image compression task, where the compressed latent features $\boldsymbol{z}_{c}$ are available.
@@
Since the residual $\boldsymbol{e}$ is unavailable during inference, we refer to DDIM [40] and assume that $\boldsymbol{z}_{n-1}$ is a linear combination of $\boldsymbol{z}_{n}$ and $\boldsymbol{z}_{0}$:
$\boldsymbol{z}_{n-1}=k_{n}\boldsymbol{z}_{0}+m_{n}\boldsymbol{z}_{n}+\sigma_{n}\epsilon.$ (5)
@@ III-C Fixed-Step Fine-Tuning Strategy
where $\lambda_{r}$ is the hyper-parameter that controls the trade-off, $R(\cdot)$ denotes the estimated rate, $sg(\cdot)$ denotes the stop-gradient operator, and $\beta=0.25$. Thanks to the proposed relay residual diffusion framework, we can achieve high-quality reconstruction in fewer than 5 denoising steps, as demonstrated in Fig. 8. This efficiency allows us to fine-tune the model using the entire reconstruction process with limited computational resources.
@@ III-D Controllable Detail Generation
Since the compressed latent feature already contains image information, directly using stable diffusion’s noise estimator $\epsilon_{sd}$ to predict noise $\epsilon_{sd}(\boldsymbol{z}_{n},n)$ results in low-frequency reconstructed images, as shown in the second column of Fig. 10. Inspired by classifier-free guidance [43], we decompose the predicted noise $\epsilon_{\theta}(\boldsymbol{z}_{n},\boldsymbol{c},n)$ into a low-frequency control component $\epsilon_{sd}(\boldsymbol{z}_{n},n)$ and a high-frequency control component $\epsilon_{\theta}(\boldsymbol{z}_{n},\boldsymbol{c},n)-\epsilon_{sd}(\boldsymbol{z}_{n},n)$, and control the balance between smoothness and sharpness by adjusting the intensity of the high-frequency control component:
$\hat{\epsilon}_{n}=\epsilon_{sd}(\boldsymbol{z}_{n},n)+\lambda_{s}(\epsilon_{\theta}(\boldsymbol{z}_{n},\boldsymbol{c},n)-\epsilon_{sd}(\boldsymbol{z}_{n},n)),$ (15)
where $\lambda_{s}$ is the guidance scale. By adjusting the value of $\lambda_{s}$, we can regulate the amount of high-frequency details introduced into the reconstructed image. In the following experiments, we set $\lambda_{s}=1$ by default unless otherwise specified.
@@ IV Experiments
Figure 4: Quantitative comparisons with state-of-the-art methods in terms of perceptual metrics (DISTS $\downarrow$ / NIQE $\downarrow$ / LPIPS $\downarrow$ / FID $\downarrow$ / KID $\downarrow$ ) on the Kodak, Tecnick ,and CLIC2020 datasets. Solid lines are used for diffusion-based methods, while dashed lines represent other methods. For RDEIC, we use 2 denoising steps for the two models with larger bpp and 5 steps for the remaining models.
@@ IV-A1 Datasets
The proposed RDEIC is trained on the LSDIR [44] dataset, which contains 84,911 high-quality images. For evaluation, we use three common benchmark datasets, i.e., the Kodak [45] dataset with 24 natural images of 768$\times$512 pixels, the Tecnick [46] dataset with 140 images of 1200$\times$1200 pixels, and the CLIC2020 [47] dataset with 428 high-quality images. For the Tecnick and CLIC2020 datasets, we resize the images so that the shorter dimension is equal to 768 and then center-crop them with 768$\times$768 spatial resolution [11].

### v3

@@ Abstract
Diffusion-based extreme image compression methods have achieved impressive performance at extremely low bitrates. However, constrained by the iterative denoising process that starts from pure noise, these methods are limited in both fidelity and efficiency. To address these two issues, we present Relay Residual Diffusion Extreme Image Compression (RDEIC), which leverages compressed feature initialization and residual diffusion. Specifically, we first use the compressed latent features of the image with added noise, instead of pure noise, as the starting point to eliminate the unnecessary initial stages of the denoising process. Second, we directly derive a novel residual diffusion equation from Stable Diffusion’s original diffusion equation that reconstructs the raw image by iteratively removing the added noise and the residual between the compressed and target latent features. In this way, we effectively combine the efficiency of residual diffusion with the powerful generative capability of Stable Diffusion. Third, we propose a fixed-step fine-tuning strategy to eliminate the discrepancy between the training and inference phases, thereby further improving the reconstruction quality. Extensive experiments demonstrate that the proposed RDEIC achieves state-of-the-art visual quality and outperforms existing diffusion-based extreme image compression methods in both fidelity and efficiency. The source code and pre-trained models are available at https://github.com/huai-chang/RDEIC.
@@ I Introduction
Figure 1: The architecture of relay diffusion and our relay residual diffusion. Relay diffusion consists of two sequential stages. In the first stage, a low-resolution latent (or image) $\boldsymbol{z}_{N+1}$ is generated from pure noise $\boldsymbol{z}_{T}$ through a standard diffusion process. This intermediate result is then upsampled and injected with noise to produce $\boldsymbol{z}_{N}$, which serves as the starting point for the super-resolution stage. This design enables the model to preserve structural consistency across resolutions while avoiding redundant sampling from pure noise in later stages, improving both quality and efficiency. Inspired by relay diffusion, our relay residual diffusion reinterpret image compression and reconstruction as an analogous pair. We replace the generation stage with a compression stage, where the input image is encoded into compressed latent features $\boldsymbol{z}_{c}$. We then inject noise into $\boldsymbol{z}_{c}$ to obtain the starting point $\boldsymbol{z}_{N}$ for reconstruction stage. Since $\boldsymbol{z}_{N}$ contains both injected noise and the residual between $\boldsymbol{z}_{c}$ and the target latent $\boldsymbol{z}_{0}$, we propose a novel residual diffusion process that jointly removes both components.
@@
In this work, we propose Relay Residual Diffusion Extreme Image Compression (RDEIC) to overcome the three limitations mentioned above. To overcome the first two limitations, we proposed a novel relay residual diffusion framework, as shown in Fig. 1(b). Specifically, we construct the starting point using the compressed latent features combined with slight noise, transitioning between the starting point and target latent features by shifting the residual between them. This approach significantly reduces the number of denoising steps required for reconstruction while ensures that the starting point retains most of the information from the compressed features, providing a strong foundation for subsequent detail generation. To leverage the robust generative capability of pre-trained stable diffusion for extreme image compression, we derive a novel residual diffusion equation directly from stable diffusion’s diffusion equation, rather than designing a new diffusion equation from scratch as done in [17]. To address the third limitation, we introduce a fixed-step fine-tuning strategy to eliminate the discrepancy between the training and inference phases. By fine-tuning RDEIC throughout the entire reconstruction process, we further improve the reconstruction quality. Moreover, to meet users’ diverse requirements, we introduce a controllable detail generation method that achieves a trade-off between smoothness and sharpness by adjusting the intensity of high-frequency components in the reconstructions. As shown in Fig. 2, the proposed RDEIC achieves state-of-the-art perceptual performance at extremely low bitrates, and significantly outperforms existing diffusion-based extreme image compression methods with fewer inference steps.
@@
1) We propose a Relay Residual Diffusion (RRD) process that effectively combines the efficiency of residual diffusion with the powerful generative capability of stable diffusion. To the best of our knowledge, we are the first to successfully integrate stable diffusion into a residual diffusion framework.
2) To eliminate the discrepancy between the training and inference phases, we design a Fixed-Step Fine-Tuning (FSFT) strategy that refines the model through the entire reconstruction process, further improving reconstruction quality.
@@ II-A Learned Image Compression
As a pioneer work, Ballé et al. [18] proposed an VAE-based image compression framework to jointly optimize the rate-distortion performance. In [19], they later introduced a hyperprior to reduce spatial dependencies in the latent representation, greatly enhancing performance. Subsequent works further improved compression models by developing various nonlinear transforms [4, 6, 7, 20] and entropy models [21, 22, 23, 24, 25]. However, optimization for rate-distortion alone often results in unrealistic reconstructions at low bitrates, typically manifested as blurring or over-smoothness [26].
To improve perceptual quality, generative models have been integrated into compression methods. Agustsson et al. [8] added an adversarial loss for lost details generation. Mentzer et al. [9] explored the generator and discriminator architectures, as well as training strategies for perceptual image compression. In [10], Muckley et al. introduced a local adversarial discriminator to enhance statistical fidelity. With the advancement of diffusion models, some efforts have been made to apply diffusion models to image compression. For instance, Yang et al. [11] innovatively introduced a conditional diffusion model as decoder for image compression. Kuang et al. [27] proposed a consistency guidance architecture to guide the diffusion model in stably reconstructing high-quality images. In addition to natural image compression, learned methods have been extended to specialized image modalities. For example, stereo image compression jointly encodes left-right image pairs by exploiting inter-view redundancy, achieving higher compression efficiency compared to independent coding [28, 29]. Meanwhile, satellite image compression focuses on efficiently handling ultra-high-resolution aerial or satellite imagery with large spatial dimensions and complex textures [30].
In our work, we adopt a VAE-based compression backbone to effectively extract and compress image information, and integrate a diffusion-based reconstruction module on the decoder side to improve reconstruction quality at extremely low bitrates.
@@ II-C Relay Diffusion
To achieve high-resolution image generation, cascaded diffusion methods [38, 39] decompose the image generation into multiple stages, with each stage responsible for super-resolution conditioning on the previous one. However, these methods still require complete resampling at each stage, leading to inefficiencies and potential mismatches among different resolutions.
Relay diffusion, as proposed by Teng et al. [40], extends the cascaded diffusion by continuing the diffusion process directly from the low-resolution output rather than restarting from pure noise, which allows the higher-resolution stages to correct artifacts from earlier stages, as shown in Fig. 1(a). This design is particularly well-suited for tasks such as image restoration and image compression, where degraded images or features are available. For instance, PASD [41] and SeeSR [42] directly embed the LR latent into the initial random noise during the inference process to alleviate the inconsistency between training and inference. ResShift [17] further constructs a Markov chain that transfers between degraded and target features by shifting the residual between them, substantially improving the transition efficiency. However, its redesigned diffusion equation and noise schedule prevent it from leveraging the robust generative capability of pre-trained stable diffusion.
@@ III Methodology
Figure 3: The proposed RDEIC. We first map a raw image $\boldsymbol{x}$ into the latent space using the encoder $\mathcal{E}$ and then perform end-to-end lossy compression to get intermediate representations $\boldsymbol{c}$ (with 256 channels) and compressed latent features $\boldsymbol{z}_{c}$. We then use $\boldsymbol{z}_{c}$ with added noise as the starting point and apply a denoising process to reconstruct the noise-free latent feature $\boldsymbol{z}_{0}$. Note that, since $\boldsymbol{c}$ can carry more information, we choose it—rather than the more compact $z_{c}$—as the input to the control module. The decoder $\mathcal{D}$ maps $\boldsymbol{z}_{0}$ back to the pixel space, to get the reconstructed image $\hat{\boldsymbol{x}}$. (a) Vanilla diffusion framework that starts from pure noise. (b) The proposed relay residual diffusion framework that starts from compressed latent features with added noise.
@@ III-A Overall Framework
Fig. 3 shows an overview of the proposed RDEIC network. We first use an encoder $\mathcal{E}$ and analysis transform $g_{a}$ to convert the input image $\boldsymbol{x}$ to its latent representation $\boldsymbol{y}$. Then we perform hyper transform coding on $\boldsymbol{y}$ with the categorical hyper model [43] and use the space-channel context model $\mathcal{C}_{m}$ to predict the entropy parameters $(\boldsymbol{\mu},\boldsymbol{\sigma})$ to estimate the distribution of quantized latent representation $\boldsymbol{\hat{y}}$ [20]. The side information $\boldsymbol{l}_{p}$ is quantized through vector-quantization, i.e., $\hat{\boldsymbol{l}}_{p}$ is the mapping of $\boldsymbol{l}_{p}$ to its closest codebook entry. Subsequently, the synthesis transform $g_{s}$ is used to obtain the intermediate representation $\boldsymbol{c}$ (with 256 channels) and the compressed features $\boldsymbol{z}_{c}$ (with 4 channels). Random noise is then added to $\boldsymbol{z}_{c}$, which is the starting point for reconstructing the noise-free latent features $\boldsymbol{z}_{0}$ through an iterative denoising process. The denoising process is implemented by a frozen pre-trained noise estimator $\epsilon_{sd}$ of stable diffusion with trainable control module for intermediate feature modulation. Note that, since $\boldsymbol{c}$ can carry more information, we choose it, rather than the more compact $\boldsymbol{z}_{c}$, as the input to control module. Finally, the reconstructed image $\boldsymbol{\hat{x}}$ is decoded from $\boldsymbol{z}_{0}$ using the decoder $\mathcal{D}$.
@@ III-B Relay Residual Diffusion
where $\epsilon_{t}\sim\mathcal{N}(0,\boldsymbol{I})$, $\alpha_{t}=1-\beta_{t}$ and $\bar{\alpha}_{t}=\prod_{i=1}^{t}\alpha_{i}$. When $t$ is large enough, the noisy latent $\boldsymbol{z}_{t}$ is nearly a standard Gaussian distribution. In practice, $T$ is typically very large, e.g., 1000, and pure noise is set as the starting point for the reverse diffusion process. The reverse process at time step t is defined as:
$z_{t-1}=\underbrace{\frac{\sqrt{\bar{\alpha}_{t-1}}\beta_{t}}{1-\bar{\alpha}_{t}}z_{0}+\frac{\sqrt{\bar{\alpha}_{t}}(1-\bar{\alpha}_{t-1}))}{1-\bar{\alpha}_{t}}z_{t}}_{\mu_{t}(z_{t},\ z_{0})}+\underbrace{\sqrt{\frac{1-\bar{\alpha}_{t-1}}{1-\bar{\alpha}_{t}}\beta_{t}}}_{\sigma_{t}}\epsilon,$ (2)
Where $\epsilon\sim\mathcal{N}(0,I)$. However, this approach is not optimal for the image compression task, where the compressed latent features $\boldsymbol{z}_{c}$ are available.
@@
Since the residual $\boldsymbol{e}$ is unavailable during inference, we refer to Eq. (2) and define $\boldsymbol{z}_{n-1}$ as a linear combination of $\boldsymbol{z}_{n}$ and $\boldsymbol{z}_{0}$:
$\boldsymbol{z}_{n-1}=\underbrace{k_{n}\boldsymbol{z}_{0}+m_{n}\boldsymbol{z}_{n}}_{\mu_{n}(\boldsymbol{z}_{n},\ \boldsymbol{z}_{0})}+\sigma_{n}\epsilon,$ (6)
@@ III-C Fixed-Step Fine-Tuning Strategy
where $\lambda_{r}$ is the hyper-parameter that controls the trade-off, $R(\cdot)$ denotes the estimated rate, $sg(\cdot)$ denotes the stop-gradient operator, and $\beta=0.25$. Thanks to the proposed relay residual diffusion framework, we can achieve high-quality reconstruction in fewer than 5 denoising steps. This efficiency allows us to fine-tune the model using the entire reconstruction process with limited computational resources.
@@ III-D Controllable Detail Generation
Figure 5: Quantitative comparisons with state-of-the-art methods in terms of perceptual metrics (DISTS $\downarrow$ / NIQE $\downarrow$ / LPIPS $\downarrow$ / FID $\downarrow$ / KID $\downarrow$ ) on the Kodak, Tecnick, and CLIC2020 datasets. Solid lines are used for diffusion-based methods, while dashed lines represent other methods.
In image generation tasks, Classifier-Free Guidance (CFG) [47] is often used to enhance conditional generation by combining unconditional and conditional noise predictions:
$\hat{\epsilon}=\epsilon_{uncond}+\omega\cdot(\epsilon_{cond}-\epsilon_{uncond}),$ (16)
where $\epsilon_{cond}$ and $\epsilon_{uncond}$ denote the noise predictions with and without conditioning, respectively. $\omega$ is the guidance scale that controls the influence of the conditional input. As $\omega$ increases, the conditional signal is amplified, resulting in outputs that are sharper and more faithful to the conditional input. Inspired by CFG, we adopt a similar principle in our reconstruction stage to control the trade-off between smoothness and sharpness. Specifically, we treat $\epsilon_{\text{sd}}(z_{n},n)$ as the unconditional prediction from the pre-trained Stable Diffusion denoiser, and $\epsilon_{\theta}(z_{n},c,n)$ as the conditional prediction guided by the conditional features $\boldsymbol{c}$. The final noise estimate is computed as:
$\hat{\epsilon}_{n}=\epsilon_{sd}(\boldsymbol{z}_{n},n)+\lambda_{s}\cdot(\epsilon_{\theta}(\boldsymbol{z}_{n},\boldsymbol{c},n)-\epsilon_{sd}(\boldsymbol{z}_{n},n)),$ (17)
where $\lambda_{s}$ plays the same role as the guidance scale $\omega$, allowing us to regulate the amount of high-frequency details introduced into the reconstructed image. In the following experiments, we set $\lambda_{s}=1$ by default unless otherwise specified.
@@ IV-A1 Datasets
The proposed RDEIC is trained on the LSDIR [48] dataset, which contains 84,911 high-quality images. For evaluation, we use three common benchmark datasets, i.e., the Kodak [49] dataset with 24 natural images of 768$\times$512 pixels, the Tecnick [50] dataset with 140 images of 1200$\times$1200 pixels, and the CLIC2020 [51] dataset with 428 high-quality images. For the Tecnick and CLIC2020 datasets, we resize the images so that the shorter dimension is equal to 768 and then center-crop them with 768$\times$768 spatial resolution [11]. The Kodak dataset was evaluated directly at its original resolution.